DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 19 Communications 326 3:45 pm

Realizing holonomic constraints

Richard Froese (University of British Columbia)

To what extent can a system of particles constrained to lie on some submanifold in configuration space be considered as a limit of systems of unconstrained particles moving under the influence of increasingly strong constraining potentials? I will discuss some answers to the question in both classical and quantum mechanics.